OpenStudy (anonymous):

The electron are accelrated to a speed of 2.40*10^7 in 1.8*10^-9, the force experinced by an electron ?

OpenStudy (anonymous):

speed of m/s in s

OpenStudy (anonymous):

@satellite73

OpenStudy (astrophysics):

\[F = ma = \frac{ mv }{ t }\] where the mass of an electron is \[\approx 9.11 \times 10^{-31} kg\]

OpenStudy (anonymous):

is used Ep=Ek is that correct?

OpenStudy (anonymous):

Eelectron=Eproton? or am i overthinking it ?

OpenStudy (astrophysics):

Well you need the force

OpenStudy (anonymous):

can i do hf=1/2mv^2?

OpenStudy (anonymous):

i think i am getting confused bettween force and freunecy

OpenStudy (astrophysics):

Yes, you are, what you're doing wouldn't make sense, that deals with the photoelectric effect.

OpenStudy (astrophysics):

You want the force F, so we use Newton's second law of motion \[F = ma\] where acceleration is \[a = \frac{ \Delta v }{ \Delta t }\]

OpenStudy (astrophysics):

So for our case we have \[F = ma = \frac{ mv }{ t }\]

OpenStudy (anonymous):

btw do u know the component method when doing momentum questions?

OpenStudy (astrophysics):

As in x and y directions

OpenStudy (astrophysics):

But yes, I know momentum, and how to deal with the problems.

OpenStudy (anonymous):

ok hold one i will get more probs

OpenStudy (anonymous):

i know E=Ef-Ei

OpenStudy (astrophysics):

\[n_m - n_2~~~m=(3,4,5)\] so if you do \[n_3 - n_1 = 3.4-1.51 = 1.89 eV\] if you do the others they would require an energy transfer above 2.0.

OpenStudy (anonymous):

in a glass vacuum electrons are accelatred trough a large potential difference

OpenStudy (anonymous):

the electrons are accelreated through a potential difference of 615 V and the spacing between the toms in the metal target is 0.243nm

OpenStudy (anonymous):

sorry 0.234nm

OpenStudy (anonymous):

@Astrophysics

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I agree with @Astrophysics

OpenStudy (anonymous):

no i am done that question i am talkign about the new one

OpenStudy (anonymous):

like from in glass vaccum

OpenStudy (michele_laino):

I see only initial data, what quantity is requested to compute, please?

OpenStudy (anonymous):

The angle of first order maximum is ?

OpenStudy (michele_laino):

we have to apply teh Bragg's formula: \[n\lambda = 2d\sin \theta \]

OpenStudy (michele_laino):

where n=1

OpenStudy (anonymous):

idk about braggg but we did the compton effect

OpenStudy (michele_laino):

\lambda is given by the De Broglie relationship: \[\Large \lambda = \frac{h}{p}\]

OpenStudy (michele_laino):

the Compton effect is related to a collision between a photon and a free electron

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

what to do next

OpenStudy (michele_laino):

I think that your exercise is related to the diffraction by electrons

OpenStudy (michele_laino):

the energy of the electrons, is: \[\Large E = \frac{{{p^2}}}{{2{m_e}}}\]

OpenStudy (anonymous):

^?

OpenStudy (michele_laino):

since our electron are classical (not relativistic) electrons. So we can write: \[\Large p = \sqrt {2{m_e}E} \]

OpenStudy (michele_laino):

please note that the energy of our electron is all kinetic energy: \[\Large E = \frac{1}{2}{m_e}{v^2} = \frac{1}{2}{m_e}{\left( {\frac{p}{{{m_e}}}} \right)^2} = \frac{{{p^2}}}{{2{m_e}}}\]

OpenStudy (anonymous):

oh r u doing somthing like Ep=Ek

OpenStudy (michele_laino):

no, I have changed variable, I expressed the kinetic energy of electron as a function of its momentum: \[\Large {v_e} = \frac{p}{{{m_e}}}\]

OpenStudy (michele_laino):

it is a simple substitution

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

now, we have to compute the momentum of our electron, as follows: \[\Large p = \sqrt {2{m_e}E} \] with your initial data. What do you get?

OpenStudy (michele_laino):

sorry after the acceleration, the kinetic energy gained by our electron is coming from the potential energy due to the electric potential, so we can write: \[\Large E = eV = ...joules\]

OpenStudy (anonymous):

2.188*10^-9

OpenStudy (michele_laino):

I got this: \[\Large E = eV = 1.6 \times {10^{ - 19}} \times 615 = 9.84 \times {10^{ - 17}}joules\]

OpenStudy (anonymous):

angle?

OpenStudy (michele_laino):

we have to compute the wavelength associated to our electron first, in order to do that, we have to compute its momentum, so we have to compute this: \[\large p = \sqrt {2{m_e}E} = \sqrt {2 \times 9.11 \times {{10}^{ - 31}} \times 9.84 \times {{10}^{ - 17}}} = ...?\]

OpenStudy (michele_laino):

please complete, what do you get?

OpenStudy (anonymous):

1.33*10^-23

OpenStudy (michele_laino):

coorrect!

OpenStudy (michele_laino):

correct!*

OpenStudy (anonymous):

angle?

OpenStudy (michele_laino):

now we can compute \lambda as follows: \[\Large \lambda = \frac{h}{p} = \frac{{6.62 \times {{10}^{ - 34}}}}{{1.34 \times {{10}^{ - 23}}}} = ...cm\]

OpenStudy (anonymous):

4.94*10^-11

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

angle?

OpenStudy (michele_laino):

finally, we can find your angle, as follows: \[\Large \sin \theta = \frac{\lambda }{{2d}} = \frac{{4.9410 - 11}}{{2 \times 2.34 \times {{10}^{ - 8}}}} = ...\]

OpenStudy (anonymous):

0.06049

OpenStudy (michele_laino):

sorry, our wavelength is measured in meters not in cm, so we have: \[\Large \sin \theta = \frac{\lambda }{{2d}} = \frac{{4.94 \times {{10}^{ - 11}}}}{{2 \times 2.34 \times {{10}^{ - 10}}}} = ...\]

OpenStudy (anonymous):

6.05

OpenStudy (michele_laino):

I got this: \[\Large \sin \theta = \frac{\lambda }{{2d}} = \frac{{4.94 \times {{10}^{ - 11}}}}{{2 \times 2.34 \times {{10}^{ - 10}}}} = 0.105555\]

OpenStudy (anonymous):

i am in the end

OpenStudy (anonymous):

from this the angle is 6.04 but the answer is 122

OpenStudy (michele_laino):

yes! we get: \[\theta = 6.06\;{\text{degrees}}\]

OpenStudy (michele_laino):

122 degrees?

OpenStudy (michele_laino):

please wait I check my computations

OpenStudy (michele_laino):

I got the same result

OpenStudy (michele_laino):

I made the same computation using a slight different procedure, and I got the same result again, in particular, the wavelength of your electrons is: \[\lambda {\text{ = 0}}{\text{.494}}\;{\text{Angstrom}}\]